OpenStudy (anonymous):

Evaluate the polynomial function at the given value. Find P(-5) if P(x) = 2x2 - 9x - 6.

OpenStudy (anonymous):

just replace all of the xs with -5 and solve

OpenStudy (anonymous):

89

OpenStudy (anonymous):

P(X)=2*2-9(-5)-6=4+45-6=43